UT 24 Aug 2020 - Night #3 of Restart

OSA: D.G. Huerta

ISp: E. Solheid

Mtn Manager: R. Hansen

ISA: B. Rothberg (Tucson - secret bunker)

Special Guest star: O. Kuhn (Tucson)

Problems

  1. All-Sky camera has been flaky all night, stopped working for a few hours, worked less than 1 hour, currently not working as of (UT 05:29)
  2. Weather station has been offline line all night - relying on Humidity and other info from SMT - fixed at UT 05:46

Night Log:

UT 01:15 - logged in will bring up MODS shortly - weather on summit looks bad - dark clouds

UT 01:22 - waking up MODS

UT 01:23 - taking sieveSnap

all OK

UT 01:29 - running test preset

-all ok

UT 02:00 - still closed for clouds

UT 02:36 - D.G.Huerta noticed very hazy outside, slightly clear overhead but ring of dark clouds coming in from the north

UT 03:12 - Opening

-will start with Rotator Center for Detector - M71

UT 03:19 - weather station appears to be offline - ENVGUI shows all red an error connecting to system, explains why the humidity has been super constant these last few nights as reported on the webapges (info.tucson.lbto.org)

UT 03:37 - pointing complete, moving to collimation

UT 03:40 - working on foxus, needs manual adjsutmentse

-some fuzziness in SH spots at btm left of AGW5 - WFS reported about 0.7", and fuzziness top right of AGW6

Rotator Center for Detector

UT 03:47 - sending Active Preset to M71 field

M1B 0002 ; M1R 0002

M2B 0002; M2R 0002

Track Preset to PA=0

M1B 0003 ; M1R 0003

M2B 0003; M2R 0003

UT 03:53 - DIMM and weather station still offline

Track Preset to PA=180

M1B 0004 ; M1R 0004

M2B 0004; M2R 0004

Track Preset to PA=90

M1B 0005 ; M1R 0005

M2B 0005; M2R 0005

Track Preset to PA=270

M1B 0006 ; M1R 0006

M2B 0006; M2R 0006

Rotator ZP Check (AGW with MOS mask)

UT 04:08 moving to Stone O for MOS acquisition

M1R 0007-0008 thru-MOS and field images

M2R 0007-0008 thru-MOS amd field images

UT 04:14 - all-sky camera is NOT updating

MODS-2 results:

Computed Mask Alignment Offset:
dX = 4.773 arcsec
dY = -1.009 arcsec
dPA = -0.011 deg

MODS-1 resultss:

Computed Mask Alignment Offset:
dX = -0.330 arcsec
dY = 0.329 arcsec
dPA = -0.157 deg

Spectrophotometric Star - all modes

UT 04:27 - acquiring on BD+28 4211

M1R

Computed Slit Alignment Offset:
dX = -0.096 arcsec
dY = 11.370 arcsec

M2R

Computed Slit Alignment Offset:
dX = 3.946 arcsec
dY = 7.224 arcsec

M1R 0010-0012 - alignment need to refine

Computed Slit Alignment Offset:
dX = -0.670 arcsec
dY = -0.175 arcsec

M2R 0010-0012 -alingment need to refine

Computed Slit Alignment Offset:
dX = -0.476 arcsec
dY = -0.335 arcsec

UT 04:33 - starting spectroscopy for both MODS

M1B 0007-0009 - Dual Grating ; M1R 0013-0015 - Dual Grating

M2B 0007-0009 - Dual Grating ; M2R 0013-0015 - Dual Grating

M1R 0016-0018 - Red Grating

M2R 0016-0018 - Red Grating

M1B 0010-0012 - Blue Grating

M2B 0010-0012 - Blue Grating

M1B 0013-0015 ; M1R 0019-0021 - Dual Prism

M2B 0013-0015; M2R 0019-0021 - Dual Prism

M1R 0022-0023 - Red Prism - M1R 23 don't use

M2R 16 is don't use

UT 04:55 - All-Sky is back up with real-time images

UT 05:14 - lost guidestar (5min ago) paused at Blue Prism M2 (1st exposure) , Red Prim M1 (1st exposure)

UT 05:18 - clsoing for clouds

UT 05:19 - all-sky camera AGAIN has stopped updating.

UT 08:32 - brought up all-sky from weather computer via OBS-2. Dark and stormy clouds! Still not updating on info website or overhead weather display.

UT 09:54 - still closed

UT 11:22 - Astronomical Twilight - still thick clouds - putting MODS to sleep, calling it for tonight.


PROPOSED PLAN OF ACTION:

  • Telescope Wake-up ~0.5hr
    • DONE test presets (sidereal and NS)
  • MODS
    • Basic Checkout
      • DONE Verify collimation
      • Verify AGW 5/6 Hotspots are centered - NOT NEEDED
      • DONE Detector Rotator Center
      • DONE Rotator ZP verification
      • AGW Transform if deemed necessary (i.e. Rot ZPs significantly changed or dithering offsets found to be problematic) - NOT NEEDED
      • Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode. -Almost completed
      • Dice-5 imaging offsets
    • Test Science OB
      • Sidereal
      • Non-Sidereal
  • LBC- Red - time permitting could coliimate, take test images on-sky

Closed Dome Activities

  • MODS Prism Calibrations
    • Dual
    • Red
    • Blue
-- BarryRothberg - 23 Aug 2020
Topic revision: r7 - 25 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback